Determine the molar mass of a gas if its pressure is to fall to one-half of its value in a vertical distance of
one meter at 298 K.

Grade:12

1 Answers

Arun
5 years ago
 

ρg(h1-h2)= P-P/2

ρ=PM/RT

h1-h2=1

R=8.314

g=9.8

T= 298

M=RT/2g=8.314*298/2*9.8

M=126.4

 

Regards

Arun (askIITians forum expert)